OpenStudy (anonymous):

when expanding ln(4x^(5)-x-1), is it correct to expand to 5ln4x-lnx-ln1?

OpenStudy (anonymous):

No, according to log rules: \[\log(a) - \log(b) = \log(a/b) , b \neq 0\] So youd need a division inside the log in order to split it into subtractions.

OpenStudy (anonymous):

thanks